# ANUBHUTI: A Comprehensive Corpus for Sentiment Analysis in Bangla Regional Languages

### v1

@@ Abstract
Sentiment analysis for regional dialects of Bangla remains an underexplored area due to linguistic diversity and limited annotated data. This paper introduces ANUBHUTI, a comprehensive dataset consisting of 2,000 sentences manually translated from standard Bangla into four major regional dialects—Mymensingh, Noakhali, Sylhet, and Chittagong. The dataset predominantly features political and religious content, reflecting the contemporary socio-political landscape of Bangladesh, alongside neutral texts to maintain balance. Each sentence is annotated using a dual annotation scheme: (i) multiclass thematic labeling categorizes sentences as Political, Religious, or Neutral, and (ii) multilabel emotion annotation assigns one or more emotions from Anger, Contempt, Disgust, Enjoyment, Fear, Sadness, and Surprise. Expert native translators conducted the translation and annotation, with quality assurance performed via Cohen’s Kappa inter-annotator agreement, achieving strong consistency across dialects. The dataset was further refined through systematic checks for missing data, anomalies, and inconsistencies. ANUBHUTI fills a critical gap in resources for sentiment analysis in low-resource Bangla dialects, enabling more accurate and context-aware natural language processing.
Keywords Regional Dialects  $\cdot$ Sentiment Analysis  $\cdot$ Low Resource Language  $\cdot$ Bangla BERT
Specification Table
Subject Computer Science Applications
Specific subject area Sentiment Analysis on Bangla Regional text data.
Type of data Table (CSV)
Data collection ANUBHUTI comprises approximately 500 sentences per region. The raw sentences were collected from a publicly available Bangla Sentiment Analysis corpus and further manual transaltion was done via human annotators. The texts primarily cover religious and political content.
Data source location MONOVAB
Data accessibility Repository Name: Mendeley Data (10.17632/gbkszkt8z3.2)
Related research article None
1 Value of the data
  • ANUBHUTI is a pioneering corpus developed for sentiment analysis in Bangla regional dialects, addressing a critical gap in low-resource NLP research for languages such as Sylhet, Chittagong, Noakhali, and Mymensingh.
  • The dataset is thoroughly annotated by expert linguists specializing in regional Bangla dialects, ensuring high-quality sentiment analysis dataset.
  • The corpus provides a foundation for practical applications such as dialect-aware chatbots, mental health analysis, and social media monitoring, contributing to inclusive and regionally adaptive AI technologies.
  • Provided in a structured CSV format, the dataset is compatible with various NLP pipelines, making it a valuable resource for future academic and industry research in Bangla language processing.
@@ 2 Related Works
Sentiment and emotion analysis in morphologically rich and low-resource languages has witnessed considerable growth in recent years. However, the majority of progress has been concentrated around high-resource languages or standardized variants, leaving dialectal and regionally diverse linguistic forms underexplored. This literature review synthesizes recent studies across Bangla, Arabic, and Hindi, with a particular emphasis on dialect-aware sentiment classification, resource development, and methodological innovation.
@@ 2.1 Bangla Sentiment and Emotion Analysis
Bangla, as a low-resource yet widely spoken language, has increasingly become the focus of sentiment and emotion analysis efforts. Bhowmik et al. (2021)[1] proposed a supervised learning framework combining a categorical weighted lexicon (LDD) with a rule-based sentiment scoring method called Bangla Text Sentiment Score (BTSC). Applied to cricket and restaurant review datasets, their SVM model achieved a classification accuracy of 82.21% using BiGram features. However, the study underscored the shallow progress of Bangla sentiment analysis due to a lack of computational resources and annotated corpora.
To address the need for large-scale data, Kabir et al. (2023)[2] introduced BanglaBook, a dataset of over 158,000 Bangla book reviews categorized into positive, neutral, and negative sentiments. The authors evaluated multiple traditional and deep learning models, including fine-tuned Bangla-BERT, which achieved a weighted F1-score of 0.933. The work emphasized the shortage of domain-specific annotated datasets and advanced pre-trained models tailored for Bangla.
In a machine-learning-focused study, Paul et al. (2025)[3] analyzed Bangla social media comments using ML classifiers, BiLSTM, and LIME for explainability. Leveraging the EmoNoBa dataset with six emotion classes, their Decision Tree model boosted with AdaBoost achieved an F1-score of 0.786. The study addressed the concern that social media–based datasets might not reflect the full spectrum of Bangla linguistic usage.
In another recent study, Mahmud et al.(2024)[4] presented a benchmark dataset for sentiment analysis of Bangla cricket-related social media content, focusing on low-resource and noisy textual environments. With a Cohen’s Kappa score of 0.971 indicating strong annotation reliability, the dataset supported evaluations using deep learning models. The BiLSTM model achieved 95.2% accuracy, outperforming both LSTM and traditional RNN models. This work addressed the scarcity of annotated Bangla datasets in the sports domain.
@@
Islam et al. (2023)[5] further advanced resource development by releasing SentiGOLD, a 70,000-sample Bangla dataset spanning over 30 domains such as news, blogs, and social media. Featuring a nuanced five-class sentiment scale, SentiGOLD achieved an Inter-Annotator Agreement of 0.88. Experiments showed that BanglaBERT performed competitively both in in-domain classification and zero-shot cross-dataset evaluations, outperforming prior datasets like SentNoB. The authors addressed issues of scale, bias, and annotation quality prevalent in earlier Bangla resources.
Beyond sentiment, Faisal et al. (2024)[6] explored emotion classification in their Bengali & Banglish dataset, which contains over 80,000 samples annotated across six emotional states. By combining data from EmoNoBa, UBMEC, and newly scraped social content, the authors captured both standard Bangla and Romanized Bangla (Banglish). Transformer models like BanglaBERT achieved an F1-score of 71.3% for Bangla and 64.6% for Banglish, revealing the linguistic complexity introduced by code-switching.
In the e-commerce domain, Haque et al. (2018)[7] analyzed sentiment in a large-scale Amazon product review dataset of over 48,000 entries across categories like electronics and musical instruments. Using active learning for data labeling and applying SVM with TF-IDF and n-gram features, the authors achieved over 90% classification accuracy. However, they noted the challenge of acquiring gold-standard annotated corpora from commercial sources.
@@ 2.2 Sentiment Analysis in Arabic Dialects
Table 1: Summary of Sentiment Analysis Study Across Languages and Datasets
Author Dataset Approaches Accuracy Baly et al.[8] MD-ArSenTD SVM, LSTM Egypt: SVM: 60.6%, LSTM: 70.0%
UAE: SVM: 51.1%, LSTM: 63.7%
Medhaffar et al.[9] TSAC SVM, Naive Bayes, MLP MLP: 78%, SVM: 77%, BNB: 58% Faria et al.[10] Vashantor mT5, BanglaT5, mBERT, Bangla-bert-base Bangla-bert: 85.86%, mBERT: 84.36% Pandey et al.[11] Hindi Review Dataset HSWN, Rule Handling HSWN-based: 80% Bhowmik et al.[1] Bangla ABSA UniGram, BiGram UniGram: 77.91%–78.69%
BiGram: 80.58%–82.21%
Kabir et al.[2] BanglaBook SVM, LSTM, Bangla-BERT Bangla-BERT (F1): 93.31% Mahmud et al.[4] Bangla Cricket Sentiment ML, DL (LR, RF, SVM, RNNs) RNN: 52.39%, LSTM: 89.72%, BiLSTM: 95.12% Islam et al.[5] SentiGOLD LSTM, BanglaBERT Macro F1: 0.62 (5-class), 0.61 (cross-set) Faisal et al.[6] Bangla + Banglish Machine Learning, Deep Learning Bangla: 71.12%, Banglish: 64.31% Haque et al.[7] Amazon Reviews NB, SVM, RF, etc. Music: 94.02%, Phones: 93.57%, Electronics: 93.52% Paul et al.[3] EmoNoBa n-gram, Machine Learning models: Decision Tree, SVM, Random Forest, KNN Decision Tree: 78.60%
@@ 2.3 Sentiment Analysis in Hindi
2.4 Dialect Based Studies
A major enabling resource for dialectal Bangla NLP is Vashantor, proposed by Faria et al. (2023)[10]. This benchmark dataset facilitates the automated translation of regional Bangla dialects into standard Bangla. Covering a wide geographic and linguistic span—including Sylhet, Noakhali, Barishal, Chittagong, and Mymensingh dialects—Vashantor is a pivotal contribution toward dialect normalization and low-resource translation. Although not focused directly on sentiment analysis, the dataset provides foundational infrastructure to build dialect-aware NLP models and enhance downstream tasks such as emotion or sentiment classification in regional varieties of Bangla. Similar to Vashantor dataset, Sultana et al. (2025)[12] and Paul et al. (2025)[13] proposed their own version of Dialect Translation dataset. In a recent study of Regional Dialect, Paul et al. (2024)[14] explored regional dialect classification using data from the VASANTOR dataset, covering Bangla dialects like Mymensingh, Chittagong, Barishal, Noakhali, and Sylhet. The study compared BanglaBERT, GPT-3.5 Turbo, and Gemini 1.5 Pro, finding that BanglaBERT achieved the highest classification accuracy of 88.74%. This work underlines the complexity of intra-language variation and the need for dialect-sensitive models.
Table 1 summarizes the literature reviews of our study. Despite significant advancements in sentiment and emotion analysis for standard Bangla, Hindi, and Arabic dialects, a notable research gap persists in the area of regional dialect sentiment analysis, particularly for low-resource languages like Bangla. Existing studies either rely heavily on standard language forms or lack the granularity to capture regional linguistic nuances. While efforts like Vashantor and regional classification models have laid the groundwork, sentiment-specific datasets and models for Bangla dialects remain virtually nonexistent. This underscores the urgent need to develop sentiment analysis systems that can understand and process non-standard dialects, which are prevalent in real-world communication but poorly represented in current NLP infrastructure. In response to this gap, our research aims to advance dialect-aware sentiment analysis by leveraging modern transformer architectures, domain adaptation, and interpretability techniques—thus contributing to a more inclusive and linguistically representative natural language processing landscape.
@@ 3 Data Description
Sentiment analysis, also known as opinion mining, is a fundamental task in Natural Language Processing (NLP) that involves identifying and categorizing emotions or opinions expressed in text. It plays a vital role in a wide range of applications, including social media analysis, political forecasting, and consumer behavior research, by assessing the polarity of a given textual input.
Although sentiment analysis has achieved substantial progress in high-resource languages like English, its development in low-resource languages such as Bangla is still in the early stages. Most existing works in Bangla sentiment analysis focus primarily on formal or standard language forms, overlooking the rich diversity of dialectal variations that exist across different regions of Bangladesh. However, dialectal expressions are often more emotionally charged, contextually grounded, and reflective of grassroots-level sentiment. Addressing this gap, our work introduces a new perspective by focusing on Bangla regional dialects for sentiment analysis. This approach enables a more inclusive and granular understanding of public opinion, especially in socio-political and religious contexts that dominate online discourse in Bangladesh.
@@
Figure 1: A Systematic Pipeline for the development of ANUBHUTI Dataset
@@ 3.1 Data Collection
For the initial phase of this research, a total of 500 raw data samples were collected to support preliminary sentiment analysis. The primary focus was on political and religious discourse, as these domains tend to evoke strong emotional reactions and diverse public opinions. A subset of neutral content was also included to provide a baseline for comparative sentiment evaluation. The choice of political and religious content was motivated by the current socio-political climate in Bangladesh, where heightened public discourse around governance, ideological tensions, and religious sentiments is prominently visible across digital platforms. These topics often reflect deeper emotional undercurrents and are highly relevant for training sentiment classification models that aim to capture context-sensitive regional expressions.
@@ 3.2 Data Translation
To construct a regionally diverse sentiment dataset, we selected four prominent dialect regions of Bangladesh: Mymensingh, Noakhali, Sylhet, and Chittagong. These regions were chosen based on their distinct linguistic features and strong presence in local digital discourse. To ensure balanced representation, we translated 500 sentences into the corresponding dialect for each region, resulting in a total of 2,000 dialectal samples.
Given the complexity and variability of regional expressions, two native dialect experts were hired for each region to carry out the translation task. These experts possess both linguistic familiarity and contextual understanding of local expressions, allowing them to preserve the semantic integrity and sentiment polarity of each sentence during translation. The collaborative process between paired experts ensured consistency, resolved ambiguities, and maintained dialectal authenticity. This translation step was crucial for adapting sentiment-labeled content into natural, expressive dialectal forms that reflect real-world language use. The translated data serves as a foundation for training sentiment analysis models that can operate effectively on informal and region-specific Bangla text.
3.3 Data Translators Identity
To ensure the linguistic and contextual accuracy of dialectal translations, we recruited eight native speakers with proven expertise in their respective regional dialects—two for each of the selected regions: Mymensingh, Noakhali, Sylhet, and Chittagong. All translators were either language professionals, postgraduate students in linguistics, or individuals with prior experience in language documentation or translation work. Their strong familiarity with regional speech patterns, idiomatic expressions, and sentiment articulation made them highly suitable for the task. A thorough screening process was conducted to assess their dialectal proficiency and ensure consistency in translation quality. Detailed information about the translators is presented in Table 2.
Table 2: Background and Expertise of Dialect Translators
@@ 3.4 Cohen’s Kappa for Translation Agreement
To ensure the quality and reliability of dialectal translations in the ANUBHUTI dataset, it was important to evaluate the inter-annotator agreement between the two translators assigned to each region. Given the subjective nature of language translation, particularly in dialects with no standardized written form, a robust agreement metric was essential.
For this purpose, Cohen’s Kappa ($\kappa$) was selected as a statistical measure to quantify the level of agreement between two raters beyond chance. Unlike simple percentage agreement, Cohen’s Kappa accounts for the probability of agreement occurring by chance, making it a more reliable metric for evaluating subjective tasks such as translation validation. A high $\kappa$ score indicates strong consistency and shared understanding between the translators, which is critical for maintaining sentiment and semantic fidelity in the dataset. As shown in Table 3, the scores for all four regions—Mymensingh, Noakhali, Sylhet, and Chittagong—fall within the range of 0.76 to 0.84, indicating substantial to almost perfect agreement.
The computation was based on a comparison between the annotations provided by the two translators per region, using the standard Cohen’s Kappa formula. For each sentence, agreement was recorded if both annotators provided semantically and sentimentally aligned translations. The observed agreement ($p_{o}$) and the expected agreement by chance ($p_{e}$) were computed from the annotation distribution, and Cohen’s Kappa ($\kappa$) was derived using:
@@
Table 3: Cohen’s Kappa Scores for Translation Agreement by Region
@@
3.5 Data Inconsistency Checking
In the context of this research on sentiment analysis using regional dialects of Bangla, ensuring consistency and quality across all translated data is a crucial step. Since the original dataset was composed in standard Bangla and manually translated into four regional dialects—Mymensingh, Noakhali, Sylhet, and Chittagong—the risk of inconsistency was inherently high due to linguistic variability, subjective translation choices, and human error. To address these challenges, a systematic and multi-stage data inconsistency checking framework was developed. This framework consists of four primary stages:
3.5.1 Re-evaluation
In this phase, all translated sentences were manually reviewed and cross-checked against the original standard Bangla texts. The goal was to ensure semantic alignment and sentiment preservation. Reviewers focused on maintaining idiomatic correctness, dialect-specific vocabulary, and the emotional tone of each sentence. This process helped guarantee that the translated dialectal texts accurately reflected the intent and polarity of the original content.

### v2

@@ Abstract
Sentiment analysis of Bangla’s regional dialects is underexplored due to linguistic diversity and a scarcity of annotated data. This paper introduces ANUBHUTI, a comprehensive dataset consisting of 10,000 sentences manually translated from standard Bangla into four major regional dialects—Mymensingh, Noakhali, Sylhet, and Chittagong. The dataset predominantly features political and religious content—reflecting Bangladesh’s contemporary socio-political landscape—alongside neutral texts for balance. Each sentence is annotated using a dual annotation scheme: (i) multiclass thematic labeling categorizes sentences as Political, Religious, or Neutral, and (ii) multilabel emotion annotation assigns one or more emotions from Anger, Contempt, Disgust, Enjoyment, Fear, Sadness, and Surprise. Expert native translators performed the translation and annotation, with quality assurance achieved via Cohen’s Kappa inter-annotator agreement, ensuring strong consistency across dialects. Systematic checks further refined the dataset for missing data, anomalies, and inconsistencies. ANUBHUTI fills a critical resource gap for sentiment analysis in low-resource Bangla dialects, enabling more accurate, context-aware natural language processing.
Keywords Regional Languages $\cdot$ Sentiment Analysis $\cdot$ Corpus $\cdot$ Dataset $\cdot$ Natural Language Processing
1 Significance of ANUBHUTI
  • ANUBHUTI is a pioneering corpus for sentiment analysis in Bangla’s regional dialects, addressing a critical gap in low-resource NLP research for dialects like Sylhet, Chittagong, Noakhali, and Mymensingh.
  • Expert linguists specializing in regional Bangla dialects thoroughly annotated the dataset, ensuring a high-quality resource for sentiment analysis.
  • The corpus facilitates dialect-aware chatbots, mental health analysis, and social media monitoring, contributing to inclusive, regionally adaptive AI technologies.
  • The dataset’s structured CSV format ensures compatibility with various NLP pipelines, making it a valuable resource for academic and industry research in Bangla language processing.
@@ 2 Related Works
Sentiment and emotion analysis in morphologically rich, low-resource languages has grown considerably in recent years. However, progress has primarily focused on high-resource languages or standardized variants, leaving dialectal and regionally diverse linguistic forms underexplored. This literature review synthesizes recent studies across Bangla, Arabic, and Hindi, with a particular emphasis on dialect-aware sentiment classification, resource development, and methodological innovation.
@@ 2.1 Bangla Sentiment and Emotion Analysis
Bangla, as a low-resource yet widely spoken language, has increasingly become the focus of sentiment and emotion analysis efforts. Bhowmik et al. (2021)[2] proposed a supervised learning framework combining a categorical weighted lexicon (LDD) with a rule-based sentiment scoring method called Bangla Text Sentiment Score (BTSC). Their SVM model achieved 82.21% classification accuracy on cricket and restaurant review datasets using BiGram features. However, the study underscored the shallow progress of Bangla sentiment analysis due to a lack of computational resources and annotated corpora. To address the need for large-scale data, Kabir et al. (2023)[5] introduced BanglaBook, a dataset of over 158,000 Bangla book reviews categorized into positive, neutral, and negative sentiments. The authors evaluated multiple traditional and deep learning models, including fine-tuned Bangla-BERT, which achieved a weighted F1-score of 0.933. The work emphasized the shortage of domain-specific annotated datasets and advanced pre-trained models tailored for Bangla.
@@
In a machine-learning-focused study, Paul et al. (2025)[11] analyzed Bangla social media comments using ML classifiers, BiLSTM, and LIME for explainability. Leveraging the EmoNoBa dataset with six emotion classes, their Decision Tree model boosted with AdaBoost achieved an F1-score of 0.786. The study addressed concerns that social media datasets may not fully reflect Bangla linguistic usage. In another recent study, Mahmud et al.(2024)[6] presented a benchmark dataset for sentiment analysis of Bangla cricket-related social media content, focusing on low-resource and noisy textual environments. With a Cohen’s Kappa score of 0.971 indicating strong annotation reliability, the dataset supported evaluations using deep learning models. The BiLSTM model achieved 95.2% accuracy, outperforming both LSTM and traditional RNN models. This work addressed the scarcity of annotated Bangla datasets in the sports domain. Islam et al. (2023)[4] further advanced resource development by releasing SentiGOLD, a 70,000-sample Bangla dataset spanning over 30 domains such as news, blogs, and social media. Featuring a nuanced five-class sentiment scale, SentiGOLD achieved an Inter-Annotator Agreement of 0.88. Experiments showed that BanglaBERT performed competitively both in in-domain classification and zero-shot cross-dataset evaluations, outperforming prior datasets like SentNoB. The authors addressed issues of scale, bias, and annotation quality prevalent in earlier Bangla resources.
@@ 2.3 Sentiment Analysis in Hindi
2.4 Dialect-Based Sentiment Analysis
A major enabling resource for dialectal Bangla NLP is Vashantor, proposed by Faria et al. (2023)[3]. This benchmark dataset facilitates the automated translation of regional Bangla dialects into standard Bangla. Covering a wide geographic and linguistic span, including Sylhet, Noakhali, Barishal, Chittagong, and Mymensingh dialects, Vashantor is a pivotal contribution to dialect normalization and low-resource translation. Although not focused directly on sentiment analysis, the dataset provides foundational infrastructure to build dialect-aware NLP models and enhance downstream tasks such as emotion or sentiment classification in regional varieties of Bangla. Similar to Vashantor dataset, Sultana et al. (2025)[12] and Paul et al. (2025)[10] proposed their own version of Dialect Translation dataset. In a recent study of Regional Dialect, Paul et al. (2024)[9] explored regional dialect classification using data from the VASANTOR dataset, covering Bangla dialects like Mymensingh, Chittagong, Barishal, Noakhali, and Sylhet. The study compared BanglaBERT, GPT-3.5 Turbo, and Gemini 1.5 Pro, finding that BanglaBERT achieved the highest classification accuracy at 88.74%. This work underlines the complexity of intra-language variation and the need for dialect-sensitive models.
Despite significant advancements in sentiment and emotion analysis for standard Bangla, Hindi, and Arabic dialects, a notable research gap persists in the area of regional dialect sentiment analysis, particularly for low-resource languages like Bangla. Existing studies either rely heavily on standard language forms or lack the granularity to capture regional linguistic nuances. While efforts like Vashantor and regional classification models have laid the groundwork, sentiment-specific datasets and models for Bangla dialects remain virtually nonexistent. This underscores the urgent need for sentiment analysis systems that understand and process non-standard dialects, prevalent in real-world communication but underrepresented in current NLP infrastructure. In response to this gap, our research aims to advance dialect-aware sentiment analysis by leveraging modern transformer architectures, domain adaptation, and interpretability techniques—thus contributing to a more inclusive and linguistically representative natural language processing landscape.
@@ 3 Data Description
Sentiment analysis, also known as opinion mining, is a fundamental task in Natural Language Processing (NLP) that involves identifying and categorizing emotions or opinions expressed in text. It is vital in applications such as social media analysis, political forecasting, and consumer behavior research, by assessing the polarity of textual input.
While sentiment analysis has progressed substantially in high-resource languages like English, its development in low-resource languages such as Bangla remains nascent. Most Bangla sentiment analysis focuses on formal or standard language, overlooking the rich diversity of dialectal variations across Bangladesh. However, dialectal expressions are often more emotionally charged, contextually grounded, and reflective of grassroots-level sentiment. Addressing this gap, our work focuses on sentiment analysis in Bangla’s regional dialects. This approach enables a more inclusive, granular understanding of public opinion, especially in Bangladesh’s socio-political and religious online discourse.
@@
Figure 1: Systematic pipeline for the development of the ANUBHUTI dataset, illustrating data collection, preprocessing, and annotation stages.
@@ 3.1 Data Collection
Initially, we collected 2,500 raw data samples for preliminary sentiment analysis. The primary focus was on political and religious discourse, as these domains tend to evoke strong emotional reactions and diverse public opinions. A subset of neutral content was also included to provide a baseline for comparative sentiment evaluation. The choice of political and religious content reflects Bangladesh’s current socio-political climate, where heightened public discourse around governance, ideological tensions, and religious sentiments is prominent on digital platforms. These topics often reflect deeper emotional undercurrents and are highly relevant for training sentiment classification models that aim to capture context-sensitive regional expressions.
@@ 3.2 Data Translation
To construct a regionally diverse sentiment dataset, we selected four prominent dialect regions of Bangladesh: Mymensingh, Noakhali, Sylhet, and Chittagong. These regions were chosen based on their distinct linguistic features and strong presence in local digital discourse. To ensure balanced representation, we translated 2,500 sentences into the corresponding dialect for each region, resulting in a total of 10,000 dialectal samples.
Due to the complexity and variability of regional expressions, we hired two native dialect experts for each region to perform the translation. These experts’ linguistic familiarity and contextual understanding of local expressions allowed them to preserve each sentence’s semantic integrity and sentiment polarity during translation. The paired experts’ collaboration ensured consistency, resolved ambiguities, and maintained dialectal authenticity. This translation was crucial for adapting sentiment-labeled content into natural, expressive dialectal forms reflecting real-world language use. The translated data serves as a foundation for training sentiment analysis models that can operate effectively on informal and region-specific Bangla text.
3.3 Data Translator Profiles
To ensure linguistic and contextual accuracy, we recruited eight native speakers with proven expertise in their regional dialects—two for each region: Mymensingh, Noakhali, Sylhet, and Chittagong. All translators were either language professionals, postgraduate students in linguistics, or individuals with prior experience in language documentation or translation work. Their familiarity with regional speech patterns, idiomatic expressions, and sentiment articulation made them suitable for the task. A thorough screening process was conducted to assess their dialectal proficiency and ensure consistency in translation quality. Detailed information about the translators is presented in Table 1.
Table 1: Background and expertise of dialect translators, including their native languages, years of experience, and familiarity with specific dialects
@@ 3.4 Cohen’s Kappa for Translation Agreement
To ensure the quality and reliability of ANUBHUTI’s dialectal translations, we evaluated inter-annotator agreement between the two translators assigned to each region. Given the subjective nature of language translation, especially in dialects without a standardized written form, a robust agreement metric was essential.
We used Cohen’s Kappa ($\kappa$) to quantify the level of agreement between the two translators, beyond chance. Unlike simple percentage agreement, Cohen’s Kappa accounts for agreement occurring by chance, making it a more reliable metric for subjective tasks like translation validation. A high $\kappa$ score indicates strong consistency and shared understanding between the translators, which is critical for maintaining sentiment and semantic fidelity in the dataset. As shown in Table 2, the scores for all four regions—Mymensingh, Noakhali, Sylhet, and Chittagong—fall within the range of 0.76 to 0.84, indicating substantial to almost perfect agreement.
The computation was based on a comparison between the annotations provided by the two translators per region, using the standard Cohen’s Kappa formula. Agreement was recorded for each sentence if both translators provided semantically and sentimentally aligned translations. The observed agreement ($p_{o}$) and the expected agreement by chance ($p_{e}$) were computed from the annotation distribution, and Cohen’s Kappa ($\kappa$) was derived using:
@@
Table 2: Cohen’s Kappa scores quantifying inter-translator agreement for dialect translation, grouped by geographic region.
@@
3.5 Data Quality Assurance
Ensuring consistency and quality across all translated data is crucial for this sentiment analysis research on Bangla’s regional dialects.Because the original dataset in standard Bangla was manually translated into four regional dialects—Mymensingh, Noakhali, Sylhet, and Chittagong—the risk of inconsistency was high due to linguistic variability, subjective translation choices, and human error. To address these challenges, a systematic and multi-stage data inconsistency checking framework was developed. This framework consists of four primary stages:
3.5.1 Semantic and Sentiment Verification
In this phase, all translated sentences were manually reviewed and cross-checked against the original standard Bangla texts. The goal was to ensure semantic alignment and sentiment preservation. Reviewers focused on idiomatic correctness, dialect-specific vocabulary, and the emotional tone of each sentence. This process helped guarantee that the translated dialectal texts accurately reflected the original content’s intent and polarity.